OpenStudy (anonymous):

HELP! Write the equation in slope-intercept form of the line that has a slope of 2 and contains the point (4,2)

OpenStudy (anonymous):

k..with the given data.... the Eqn. of a line = > y -y1 = m(x-x1) => y-2 = 2 (x-4) => y-2 = 2x - 8 =>-2x+y = -8+2 => -2x +y = -6 => 2x-y = 6 or 2x-y-6=0 >>>>ans

OpenStudy (anonymous):

Slope- Intercept form is y - y1 = m(x - x1) where m is your slope, so just substitute in what they've given you. y - 2 = 2(x - 4) y -2 = 2x - 8 y = 2x - 6

OpenStudy (anonymous):

@EmmaMoer :Thank you very much but step by step to learn !:) Thank you again !:)

OpenStudy (anonymous):

Umm? You aren't the one that asked the question?... @E.ali

OpenStudy (anonymous):

@ShortStuff did.

OpenStudy (anonymous):

Well, YOUR WELCOME! :) @E.ali

OpenStudy (anonymous):

THANKS AGAIN MODOM !

OpenStudy (anonymous):

That did not help me at all.

OpenStudy (anonymous):

May I help you ?:)

OpenStudy (anonymous):

Thanks but someone else already did.

OpenStudy (anonymous):

But I do need help on another problem.

OpenStudy (anonymous):

Sure ! Ask !:)

OpenStudy (anonymous):

Choose the equation of the line that contains the points (-3,2) and (5,-5) Answer choices are: A:y=-7/8x-5/8 B:y=-7/8x+5/8 C:y=-7/8-5/8x

OpenStudy (anonymous):

Look : Things you see before x is slope right ?:) (In equations :) )

OpenStudy (anonymous):

you need to first find the slope using the slope formula : slope(m) = (y2 - y1) / (x2 - x1) (-3,2) x1 = -3 and y1 = 2 (5,-5) x2 = 5 and y2 = -5 now lets sub our info into the formula m = (-5 - 2) / (5 - (-3) m = -7/8 now we will use this formula : y - y1 = m(x - x1) m = -7/8 (-3,2) x1 = -3 and y1 = 2 now lets sub y - 2 = -7/8(x - (-3) y - 2 = -7/8(x + 3) y - 2 = -7/8x - 21/8 y = -7/8x - 21/8 + 2 y = -7/8x -21/8 + 16/8 y = -7/8 - 5/8 any questions ?

OpenStudy (anonymous):

@ShortStuff : Got it ?

OpenStudy (anonymous):

Yeah